OpenStudy (anonymous):

There are 20 children on a play ground. 13 played kick ball and 15 played soccer and 2 did not play either sport. How many children played both kick ball and soccer?

OpenStudy (anonymous):

A)5 B)7 C)9 D)10 E)13

OpenStudy (anonymous):

 S+B = 15 K+B = 13 S+B+K+N = 20

OpenStudy (anonymous):

N=2 => S+B+K = 18 so a system with 3 equations and 3 unknowns

OpenStudy (texaschic101):

would it work out the same if you just said 18 students...and just left off the 2 that didn't play either ??

OpenStudy (anonymous):

hold on i'll try to work on it

OpenStudy (anonymous):

is it 10?

OpenStudy (anonymous):

is it?

OpenStudy (texaschic101):

I am not 100% sure...but that is what I am getting

OpenStudy (anonymous):

tell me don't ask

OpenStudy (anonymous):

I'll show my work.

OpenStudy (anonymous):

love it

OpenStudy (anonymous):